OpenStudy (anonymous):

What are the coordinates of the x-intercepts of the graph of y = 2x2 + 6x – 20?

OpenStudy (callisto):

x-int => y=0 Put y=0 into the equation. 0 = 2x^2 + 6x – 20 0 = 2 ( x^2 +3x - 10) 0 = (x+5)(x-2) x+5 =0 or x-2 =0 Can you solve x? After solving x, you can get the coordinates (x1, 0) and (x2, 0)